# Adaptation of Agentic AI

### v2

@@ Abstract
Cutting-edge agentic AI systems are built on foundation models that can be adapted to plan, reason, and interact with external tools to perform increasingly complex and specialized tasks. As these systems grow in capability and scope, adaptation becomes a central mechanism for improving performance, reliability, and generalization. In this paper, we unify the rapidly expanding research landscape into a systematic framework that spans both agent adaptations and tool adaptations. We further decompose these into tool-execution–signaled and agent-output–signaled forms of agent adaptation, as well as agent-agnostic and agent-supervised forms of tool adaptation. We demonstrate that this framework helps clarify the design space of adaptation strategies in agentic AI, makes their trade-offs explicit, and provides practical guidance for selecting or switching among strategies during system design. We then review the representative approaches in each category, analyze their strengths and limitations, and highlight key open challenges and future opportunities. Overall, this paper aims to offer a conceptual foundation and practical roadmap for researchers and practitioners seeking to build more capable, efficient, and reliable agentic AI systems.
@@
Adaptation of Agentic AI
Background (§2) Agentic AI Systems (§2.1) Adaptation (§2.2) Prompt Engineering (§2.2.1) Fine-Tuning (§2.2.2) Overview (§3) Mathematical Notations (§3.1) Adaptation Paradigms of Agentic AI (§3.2) Illustrative Examples (§3.3) Agent Adaptation (§4) A1: Tool Execution Result as Signal (§4.1) Earlier Works: SFT & Off-Policy Methods (§4.1.1) RLVR-Based Methods (§4.1.2) A2: Agent Output as Signal (§4.2) Adaptation w/o Tools (§4.2.1) Adaptation w/ Tools (§4.2.2) Tool Adaptation (§5) T1: Agent-Agnostic Tool Adaptation (§5.1) Foundational Systems and Architectures (§5.1.1) Categories and Training Methods (§5.1.2) T2: Agent-Supervised Tool Adaptation (§5.2) Earlier Methods (§5.2.1) Subagent-as-Tool (§5.2.2) Agentic Memory and Others (§5.2.3) Comparison (§6) A1 & A2 (§6.2) T1 & T2 (§6.3) Strategic Recommendations (§6.5) Applications (§7) Deep Research (§7.1) Software Development (§7.2) Computer Use (§7.3) Drug Discovery & Development (§7.4) … Opportunities (§8) Co-Adaptation (§8.1) Continual Adaptation (§8.2) Safe Adaptation (§8.3) Efficient Adaptation (§8.4)
@@ 1 Introduction
The rapid progress of foundation models, such as large language models (LLMs), has catalyzed the rise of agentic AI systems: autonomous AI systems capable of perceiving their environment, invoking external tools, managing memory, and executing multi-step plans toward completing complex tasks [1, 2, 3, 4]. Agentic AI demonstrates remarkable potential in applications ranging from scientific discovery [5, 6] to software development and clinical research [7, 8, 9]. However, current agentic AI systems still struggle with challenges such as unreliable tool use, limited long-horizon planning, domain-specific reasoning gaps, robustness issues in real-world environments, and poor generalization to unexplored environments where the agent lacks prior interaction experience [10, 11, 12, 13]. These limitations reveal that even highly capable foundation models often require additional adaptation to specialize for particular tasks or real-world scenarios. This motivates the need for adaptation in agentic AI systems, whereby the components of an agentic system are modified or optimized so that the agent achieves higher task performance, improved reliability, and better generalization across diverse scenarios.
Building on this motivation, we conduct a comprehensive survey on the adaptation in agentic AI systems, aiming to systematically analyze how components in agentic AI systems are modified to overcome current limitations. Compared with existing surveys on modern AI agents [1, 14, 15, 16, 17, 18], this paper centers specifically on adaptation in agentic AI. To structure this rapidly expanding literature, we introduce a unified framework that organizes adaptation in agentic AI into four core paradigms spanning both agent adaptation and tool adaptation, as shown in Figure 1. This framework clarifies the underlying design space, highlights the trade-offs between different adaptation strategies, and provides practical guidance for choosing or transitioning between paradigms based on supervision signals, task requirements, and system-level constraints.
In our framework, we conclude adaptation strategies for agentic AI into two dimensions according to which component is optimized (§3). The first dimension, which we term Agent Adaptation, focuses on modifying the agent’s internal parameters, representations, or behavioral policies to better align with task requirements. This includes both traditional fine-tuning approaches [19] and modern reinforcement learning methods that leverage environment feedback [20, 21]. The second dimension, Tool Adaptation, shifts the optimization target from the agent to its external tools, e.g., retrievers, planners, memory modules, and specialized models, enabling frozen agents to benefit from an adaptive operational environment [22, 11, 23]. Within these two broad paradigms, we further identify four distinct adaptation strategies, forming a comprehensive taxonomy that organizes the rapidly evolving landscape of agentic AI research:
@@
  • T1: Agent-Agnostic Tool Adaptation3.2.3, §5.1): Tools are trained independently of the frozen agent. These tools include retrievers, domain-specific models, and other pretrained components that can be used as plug-and-play modules orchestrated by the frozen agent.
@@
It is worth noting that these four strategies are not mutually exclusive: state-of-the-art systems increasingly combine multiple adaptation paradigms to achieve optimal performance [24, 25, 26]. For instance, a deep research system might employ T1-style retrieval tools (pre-trained dense retrievers), T2-style adaptive search agents (trained via frozen LLM feedback), and A1-style reasoning agents (fine-tuned with execution feedback) in a cascaded architecture [6].
@@
In §6, we further emphasize that the choice among these paradigms involves fundamental trade-offs along several dimensions. (1) Cost and flexibility: Agent adaptation (A1/A2) typically requires substantial computational resources for training billion-parameter models but offers maximal flexibility, while tool adaptation (T1/T2) optimizes external components at lower cost but may be constrained by the frozen agent’s capabilities [27, 22]. (2) Generalization: T1 tools trained on broad data distributions often generalize well across agents and tasks [23, 28], whereas A1 methods may overfit to specific environments unless carefully regularized [20]. (3) Modularity: T2 approaches enable independent tool upgrades without agent retraining [29, 22], facilitating continuous system improvement, while A1/A2 methods may suffer from catastrophic forgetting when adapting to new tasks.
Scope and contributions. This paper provides the first comprehensive taxonomy of adaptation strategies for agentic AI, systematically organizing recent advances across agent adaptation (A1, A2) and tool adaptation (T1, T2). We offer several key contributions:
  • A unified conceptual framework that clarifies the associations and distinctions between adaptation paradigms and their underlying principles (Figure 2).
  • Detailed technical surveys of representative methods within each category, documenting their training objectives, architectural choices, and empirical performance across diverse benchmarks.
  • Systematic comparison of adaptation strategies along dimensions of cost, flexibility, generalization capability, and modularity.
  • Demonstrating how adaptation strategies are tailored to domain applications, spanning deep research, software development, computer use, and drug discovery (§7).
  • Identification of open challenges and future research directions, including unified agent-tool co-adaptation frameworks, theoretical understanding of adaptation dynamics, and standardized evaluation protocols (§8).
Organization. The remainder of this paper is organized as follows. Section 2 provides foundational concepts, introducing the core components of agentic AI systems and the two primary forms of adaptation (prompt engineering and fine-tuning). Section 3 presents an overview of adaptation paradigms under our proposed framework, formalizing the four paradigms (A1, A2, T1, T2) and illustrating them with concrete examples. Sections 4 and 5 present our main taxonomy, systematically reviewing agent adaptation (A1, A2) and tool adaptation (T1, T2) methods respectively. Section 6 compares these paradigms along key dimensions. Section 7 examines real-world applications across multiple domains. Finally, Section 8 discusses open challenges and future research directions. Throughout, we emphasize the complementary nature of agent and tool adaptation, arguing that the most effective agentic systems will strategically combine both paradigms to achieve robust, efficient, and generalizable performance across diverse tasks and environments.
@@ 2 Background
In this section, we provide the background to facilitate a better understanding of the concepts discussed throughout this survey. Specifically, we first introduce the fundamental components of Agentic AI Systems2.1). We then discuss different forms of Adaptation2.2), which enable agentic AI systems to better adjust their behaviors and capabilities to specific tasks or application scenarios.
@@ 2.1 Agentic AI Systems
Agentic AI systems refer to autonomous artificial intelligence systems capable of perceiving, reasoning, acting, and continuously improving through interaction with their environment. Such systems are designed to perform complex, open-ended tasks that require adaptive decision-making, contextual understanding, and iterative problem solving. In this survey, we primarily focus on single-agent systems, which provide a controlled yet expressive framework to study how an individual agent perceives, plans, and acts within an environment. Single-agent settings serve as the foundational building blocks of more complex multi-agent systems, in which multiple agents coordinate, cooperate, or compete to achieve shared or opposing goals. Comprehensive overviews of AI agent architectures and their extensions to multi-agent scenarios can be found in recent surveys such as [1, 2, 30].
@@
At the core of an agentic AI system lies a foundation model, typically implemented as a large language model (LLM) or multimodal model that functions as the agent’s reasoning and control center. This foundation model provides the fundamental abilities for understanding, reasoning, planning, and interaction. Complementing this core are several additional components that extend the agent’s autonomy and enable it to operate effectively in complex and dynamic environments:
@@
  • Planning Module: Decomposes complex goals into actionable steps and organizes their sequential or hierarchical execution. Depending on the degree of feedback integration, planning can be conducted in two main modes. Static planning methods, such as Chain-of-Thought [31] and Tree-of-Thought [32], enable structured reasoning through single-path or multi-path task decomposition. In contrast, dynamic planning approaches, such as ReAct [33] and Reflexion [34], incorporate feedback from the environment or past actions, allowing the agent to iteratively refine its plans and improve performance in long-horizon or partially observable scenarios.
@@
  • Tool Use: Enables the agent to interact with external resources and computational systems, extending its capabilities beyond the limitations of its internal knowledge. Typical tools include web search engines, APIs, code execution environments, Model Context Protocols (MCPs), and browser automation frameworks [35, 3]. Effective tool use involves selecting appropriate tools, constructing task-specific inputs, invoking external functions, and integrating their outputs into the agent’s reasoning and decision-making process, thereby enhancing performance in real-world and computationally intensive scenarios.
@@
  • Memory Module: Allows the agent to retain, retrieve, and utilize past information for context-aware reasoning and long-term consistency. Memory is typically divided into short-term memory, which stores contextual information generated during the current task, and long-term memory, which persists across sessions to accumulate reusable knowledge and experience [1, 36]. To access relevant information from long-term memory, many systems employ retrieval-augmented generation (RAG) mechanisms that retrieve and integrate stored knowledge into the agent’s reasoning process. Designing an effective memory module involves challenges such as how to structure stored information, when and what to retain, how to retrieve relevant knowledge efficiently, and how to seamlessly integrate it into ongoing reasoning and decision-making. [^1] While memory is a fundamental component of an agent, this survey classifies adaptive memory systems under the Tool Adaptation paradigm (specifically T2, discussed in §5.2.3). We frame them as external, optimizable tools, such as retrievers or reflective databases, that are ”tuned” using the frozen agent’s outputs as supervision.
@@ 2.2 Adaptation
Adaptation is a crucial aspect of agentic AI systems, enabling them to operate effectively across diverse and complex tasks. It allows an agent to adjust its behaviors, decision strategies, and internal representations to better align with the requirements of a specific domain, task, or operational environment. Without such adaptive mechanisms, agents may struggle to generalize beyond their initial design or handle dynamic, real-world conditions. These mechanisms can be broadly categorized into prompt-based adaptation (§2.2.1) and fine-tuning-based adaptation (§2.2.2).
@@ 2.2.1 Prompt Engineering
Prompt engineering serves as a lightweight form of adaptation that guides the behavior of an agentic AI system without modifying its underlying model parameters. Instead of retraining the core model, the agent’s behavior is shaped by carefully crafted input prompts that define goals, constraints, and contextual instructions. Through prompt design, an agent can be steered toward specific reasoning patterns, task formulations, or action strategies, enabling rapid adaptation across diverse tasks and environments.
A prompt refers to the input context provided to the agent’s core model, typically consisting of instructions, examples, or task descriptions that specify the desired behavior. By modifying or composing prompts, an agent can be adapted to new goals or environments without any additional model training, making this approach highly efficient and easily transferable across tasks. Such prompt-based adaptation has been widely adopted in recent agentic systems, such as CAMEL [37], AutoGen [38], MetaGen [39] and ChatDev [40]. For a comprehensive overview of prompt engineering techniques and their design principles, we refer readers to the survey by Sahoo et al. [41].
@@ 2.2.2 Fine-Tuning
In contrast to prompt engineering, fine-tuning achieves adaptation by updating the internal parameters of the core model. Through exposure to task-specific data, fine-tuning enables the model to internalize new knowledge, reasoning patterns, or behavioral tendencies that better align with the target domain or task objectives.
Fine-tuning can be performed at different granularities depending on data availability, computational cost, and the desired degree of adaptation. Full fine-tuning updates all model parameters using labeled data, providing maximal flexibility but often requiring substantial resources. Alternatively, parameter-efficient fine-tuning (PEFT) methods, such as low-rank adaptation (LoRA) [19], update only a small subset of parameters. These approaches offer a practical balance between efficiency and performance, enabling large agentic systems to be specialized for particular tasks or environments without extensive retraining. For a comprehensive overview of PEFT methods, we refer readers to the survey by Han et al. [42].
Fine-tuning for adapting agents encompasses several major training paradigms. Supervised Fine-Tuning (SFT) [43] performs imitation learning on curated demonstrations. Preference-based methods, such as Direct Preference Optimization (DPO) [44] and its extensions [45], align the model with human or automated preference signals. Reinforcement-learning-based approaches, including algorithms such as Proximal Policy Optimization (PPO) [46] and Group Relative Policy Optimization (GRPO) [47], further adapt agents by optimizing their behavior through interaction with evaluative environments. These families of methods form the core techniques for adapting foundation-model-based agents to specialized tasks and deployment settings. For a more comprehensive review of these approaches, see the survey by Zhang et al. [48].
We next introduce a general framework that categorizes existing agentic AI adaptation approaches. This framework, presented in the following section, forms the conceptual foundation for the rest of this survey.

### v3

@@ Abstract
Large language model (LLM) agents are moving beyond prompting alone. ChatGPT marked the rise of general-purpose LLM assistants, DeepSeek showed that on-policy reinforcement learning with verifiable rewards can improve reasoning and tool use, and OpenClaw highlights a newer direction in which agents accumulate persistent memory and reusable skills. Yet the research landscape remains fragmented across post-training, retrieval, memory, and skill systems. This survey studies these developments under a single notion of adaptation: improving an agent, its tools, or their interaction after pretraining. We organize the field with a four-paradigm framework spanning agent adaptation and tool adaptation. On the agent side, A1 (tool-execution-signaled) and A2 (agent-output-signaled) improve the agent itself through supervised fine-tuning, preference optimization, and reinforcement learning with verifiable rewards. On the tool side, T1 (agent-agnostic) provides reusable pre-trained modules any agent can call, while T2 (agent-supervised) uses the agent’s outputs to train memory systems, skill libraries, or lightweight subagents. Using this framework, we review post-training methods, adaptive memory architectures, and agent skills; compare their trade-offs in cost, flexibility, and generalization; and summarize evaluation practices across deep research, software development, computer use, and drug discovery. We conclude by outlining open problems in agent-tool co-adaptation, continual learning, safety, and efficient deployment.
@@
{forest}
@@ 1 Introduction
The rapid progress of foundation models, large language models (LLMs) in particular, has enabled a new class of agentic AI systems that perceive their environment, invoke external tools, manage memory, and execute multi-step plans to complete complex tasks [1, 2, 3, 4]. These systems show strong potential across scientific discovery [5, 6], software development, and clinical research [7, 8, 9]. Yet current agentic systems remain limited by unreliable tool use, shallow long-horizon planning, domain-specific reasoning gaps, and poor generalization to environments where the agent lacks prior interaction experience [10, 11, 12, 13]. Even the most capable foundation models require additional adaptation to specialize for particular tasks or deployment scenarios.
Three mechanisms drive this adaptation process. First, post-training (supervised fine-tuning, reinforcement learning with verifiable rewards, preference optimization) modifies the agent’s parameters to improve reasoning and tool use. Second, memory systems (episodic buffers, reflective databases, structured knowledge graphs) allow agents to retain and build upon past experience without retraining the core model. Third, skills—reusable units of procedural knowledge encoding how to perform tasks [14]—accumulate through diverse mechanisms: agents internalize tool-use procedures via post-training, while external skill libraries support discovery, invocation, and refinement across sessions. These three mechanisms interact within a shared design space that this survey maps.
Unlike existing surveys on AI agents [1, 15, 16, 17, 18, 19], we focus specifically on adaptation. We introduce a unified framework that organizes agentic adaptation into four paradigms spanning both agent and tool adaptation (Figure 1), exposing trade-offs that guide paradigm selection based on supervision signals, task requirements, and system-level constraints.
Our framework organizes adaptation along two dimensions according to which component is optimized (§3). Agent Adaptation modifies the agent’s internal parameters, representations, or behavioral policies to meet task requirements, through fine-tuning [20] or reinforcement learning with environment feedback [21, 22]. Tool Adaptation shifts the optimization target to the agent’s external tools (retrievers, planners, memory modules, specialized models), enabling frozen agents to benefit from an adaptive operational environment [23, 11, 24]. Within these two dimensions, we identify four distinct adaptation strategies:
@@
  • T1: Agent-Agnostic Tool Adaptation3.2.3, §5.1): Tools are trained independently of the frozen agent. These tools include retrievers, domain-specific models, and other pre-trained components that can be used as plug-and-play modules orchestrated by the frozen agent.
@@
These four strategies overlap in practice; we treat them as a working classification. When a method simultaneously modifies both the agent and a tool, we assign it to the paradigm corresponding to its dominant locus of optimization and note the secondary component explicitly. Current leading systems combine multiple paradigms. A deep research system may employ T1-style retrieval tools (pre-trained dense retrievers), T2-style adaptive search agents (trained via frozen LLM feedback), and A1-style reasoning agents (fine-tuned with execution feedback) in a cascaded architecture [25, 26, 27, 6].
@@
The choice among these paradigms involves trade-offs along several dimensions (§6). (1) Cost and flexibility: Agent adaptation (A1/A2) requires substantial computational resources for training billion-parameter models but offers broad flexibility, while tool adaptation (T1/T2) optimizes external components at lower cost but is constrained by the frozen agent’s capabilities [28, 23]. (2) Generalization: T1 tools trained on broad data distributions generalize well across agents and tasks [24, 29], whereas A1 methods may overfit to specific environments unless carefully regularized [21]. (3) Modularity: T2 approaches enable independent tool upgrades without agent retraining [30, 23], while A1/A2 methods risk catastrophic forgetting when adapted to new tasks.
Scope and contributions. By grounding the four-paradigm taxonomy in three concrete mechanisms—post-training, memory, and skills—we connect the abstract framework to the techniques that practitioners deploy in practice. Our key contributions are:
  • A unified $2\times 2$ framework (Figures 1, 3, 7) that decomposes agentic adaptation along two axes—what is adapted (agent vs. tool) and how the signal is obtained (execution-grounded vs. output-evaluated)—yielding four paradigms (A1/A2/T1/T2) with distinct cost, flexibility, and generalization profiles, together with cross-paradigm empirical synthesis6) showing, e.g., that T2 tool adaptation can match A2 agent training accuracy with far fewer examples in retrieval settings, and that A1-trained agents can be frozen and redeployed as T1 tools.
  • A systematic treatment of agentic memory and skills as adaptation mechanisms. We show how adaptive memory systems (episodic buffers, reflective databases, knowledge graphs) instantiate T2 adaptation, and how skill libraries bridge A1/A2 post-training with T1/T2 tool ecosystems (§5).
  • A multi-dimensional evaluation framework7) that maps benchmarks to paradigms, distinguishes verifiable execution metrics from holistic utility metrics, and identifies gaps in current evaluation practice, particularly for T2 and cross-paradigm comparison.
  • Domain-specific paradigm mapping across deep research, software development, computer use, and drug discovery (§8), showing how the dominant adaptation paradigm varies with the availability of verifiable feedback and the cost of agent retraining.
  • Identification of open challenges including agent-tool co-adaptation, continual adaptation under non-stationary task distributions, safe exploration during on-policy RL, and efficient adaptation for resource-constrained deployment (§9).
Survey methodology and scope. We surveyed the literature on agentic adaptation published through early 2026, drawing primarily from top-tier venues (NeurIPS, ICML, ICLR, ACL, EMNLP, NAACL) and high-impact arXiv preprints. Methods are assigned to paradigms based on their dominant locus of optimization and signal source; boundary cases are flagged explicitly. Quantitative cross-paradigm comparisons (e.g., data efficiency of T2 vs. A2) are drawn from case studies that may differ in confounding factors; we note these limitations where they arise.
Organization. Section 2 introduces foundational concepts. Section 3 formalizes the four paradigms with illustrative examples. Sections 4 and 5 review agent adaptation (A1, A2) and tool adaptation (T1, T2) methods, respectively. Section 6 compares paradigms along key dimensions. Section 7 presents a multi-dimensional evaluation framework. Section 8 examines real-world applications with explicit paradigm mapping. Section 9 discusses open challenges, and Section 10 concludes the paper.
@@ 2 Background
This section introduces the background necessary for the remainder of the survey. We first describe the core components of Agentic AI Systems2.1), then discuss the two primary forms of Adaptation2.2) through which these systems specialize for particular tasks or deployment scenarios.
@@ 2.1 Agentic AI Systems
An agentic AI system combines a foundation model with planning, tool use, and memory modules to autonomously execute multi-step tasks, using environmental feedback to refine its behavior over time. This survey focuses on single-agent systems, which provide a controlled yet expressive setting to study how an individual agent perceives, plans, and acts. Understanding single-agent adaptation is a prerequisite for studying multi-agent systems, in which multiple agents coordinate, cooperate, or compete; see [1, 2, 31] for overviews.
@@
At the core of an agentic AI system lies a foundation model, typically a large language model (LLM) or multimodal model that serves as the agent’s reasoning and control center. Several additional components extend the agent’s autonomy:
@@
  • Planning Module: Decomposes complex goals into actionable steps and organizes their sequential or hierarchical execution. Depending on the degree of feedback integration, planning takes two forms. Static planning methods, such as Chain-of-Thought [32] and Tree-of-Thought [33], enable structured reasoning through single-path or multi-path task decomposition. In contrast, dynamic planning approaches, such as ReAct [34] and Reflexion [35], incorporate feedback from the environment or past actions, allowing the agent to iteratively refine its plans and improve performance in long-horizon or partially observable scenarios.
@@
  • Tool Use: Enables the agent to interact with external resources and computational systems, extending its capabilities beyond the limitations of its internal knowledge. Typical tools include web search engines, APIs, code execution environments, Model Context Protocols (MCPs), and browser automation frameworks [36, 3]. The tool-selection policy itself is an adaptation surface: agents must learn which tool to invoke for a given subtask and how to parse the tool’s output for downstream reasoning.
@@
  • Memory Module: Allows the agent to retain, retrieve, and utilize past information for context-aware reasoning and long-term consistency. Memory is typically divided into short-term memory, which stores contextual information generated during the current task, and long-term memory, which persists across sessions to accumulate reusable knowledge and experience [1, 37]. To access relevant information from long-term memory, many systems employ retrieval-augmented generation (RAG) mechanisms that retrieve and integrate stored knowledge into the agent’s reasoning process. Designing an effective memory module involves challenges such as how to structure stored information, when and what to retain, how to retrieve relevant knowledge efficiently, and how to integrate it into ongoing reasoning and decision-making. [^1] While memory is a fundamental component of an agent, this survey treats most adaptive memory systems under the Tool Adaptation paradigm (discussed in §5.2.3). External, non-parametric memory stores (e.g., episodic buffers, reflective databases, skill libraries) that are updated using the frozen agent’s outputs are classified as T2; pre-trained memory modules that operate independently of a specific agent fall under T1; and parametric or hybrid memory that modifies model weights occupies the boundary between tool and agent adaptation. A detailed discussion of these distinctions appears in §3.
@@ 2.2 Adaptation
Adaptation modifies the agent or its tools over time to improve performance on specific domains, tasks, or deployment conditions. Without adaptation, even capable foundation models struggle to generalize beyond their pre-training distribution. We distinguish two broad categories: prompt-based adaptation (§2.2.1) and fine-tuning-based adaptation (§2.2.2).
@@ 2.2.1 Prompt Engineering
Prompt engineering is a lightweight form of adaptation that guides the behavior of an agentic system without modifying model parameters. The agent’s behavior is shaped by carefully crafted input prompts that define goals, constraints, and contextual instructions.
A prompt refers to the input context provided to the agent’s core model, typically consisting of instructions, examples, or task descriptions that specify the desired behavior. By modifying or composing prompts, an agent can be adapted to new goals or environments without any additional model training, making this approach efficient and transferable across tasks. Such prompt-based adaptation has been widely adopted in recent agentic systems, such as CAMEL [38], AutoGen [39], MetaGPT [40] and ChatDev [41]. For a comprehensive overview of prompt engineering techniques and their design principles, we refer readers to the survey by Sahoo et al. [42].
@@ 2.2.2 Fine-Tuning
In contrast, fine-tuning adapts the agent by updating the internal parameters of the core model. Training on task-specific data shifts the model’s distribution toward domain-appropriate behavior.
Fine-tuning can be performed at different granularities depending on data availability, computational cost, and the desired degree of adaptation. Full fine-tuning updates all model parameters using labeled data, providing maximal flexibility but often requiring substantial resources. Alternatively, parameter-efficient fine-tuning (PEFT) methods, such as low-rank adaptation (LoRA) [20], update only a small subset of parameters. LoRA, for example, typically matches full fine-tuning quality while updating less than 1% of parameters, making it practical for large agentic systems. For a comprehensive overview of PEFT methods, we refer readers to the survey by Han et al. [43].
Fine-tuning for adapting agents encompasses several major training paradigms. Supervised Fine-Tuning (SFT) [44] performs imitation learning on curated demonstrations. Preference-based methods, such as Direct Preference Optimization (DPO) [45] and its extensions [46], align the model with human or automated preference signals. Reinforcement learning methods such as Proximal Policy Optimization (PPO) [47] and Group Relative Policy Optimization (GRPO) [48] optimize behavior through environment interaction; they require only a reward signal but are harder to stabilize than SFT. For a more comprehensive review of these approaches, see the survey by Zhang et al. [49].